# Cops and Robbers Pebbling in Graphs

### v3

@@ 2.2 Upper Bounds
For integers $k$ and $d$, define the theta graph ${\Theta}(k,d)$ as the union of $k$ internally disjoint $xy$-paths, each of length $d$. Then ${\Theta}={\Theta}(k,2d)$ has $n=k(d-1)+2$, $c({\Theta})=2$, ${\sf capt}_{2}({\Theta})=d$, ${\sf gir}({\Theta})=4d$, ${\gamma}({\Theta})=k\lceil(2d-3)/3\rceil$, and ${\gamma}_{d}=2$. Thus Theorem 5 yields an upper bound of roughly $34kd/$ 3, while Theorems 12 and 15 both yield the upper bound of $2^{d+1}$, which is better or worse than Theorem 5 when $k$ is bigger or less than, roughly, $3\cdot 2^{d-1}/d$.
@@
Consider the $(3,7)$-cage McGee graph $M$, defined by $V=\{v_{i}\mid i\in{\mathbb{Z}}_{24}\}$, with $v_{i}\sim v_{i+1}$ for all $i$, $v_{i}\sim v_{i+12}$ for all $i\equiv 0\pmod{3}$, and $v_{i}\sim v_{i+7}$ for all $i\equiv 1\pmod{3}$. We have ${\gamma}_{2}(M)\leq 4$ (e.g. $\{v_{0},v_{6},v_{9},v_{15}\}$), and so ${{\pi}^{*}}(M)\leq{{\pi}^{\sf c}}(M)\leq 16$ by Theorem 15. However, this bound is not tight, as ${{\pi}^{\sf c}}(M)\leq 12$: the vertex set $\{v_{i}\mid i\equiv 0\pmod{3}\}$ induces a matching of size 4 — for each edge, place 2 cops on one of its vertices and 1 cop on the other. Incidentally, this yields ${{\pi}^{*}}(M)\leq 12$; the best known lower bound on ${{\pi}^{*}}$ comes from Result 16: ${{\pi}^{*}}(M)\geq\lceil{\hat{\pi}^{*}}(M)\rceil=\lceil 64/7\rceil=10$. Hence we are left with a gap in the bounds for $M$: $10\leq{{\pi}^{*}}(M)\leq{{\pi}^{\sf c}}(M)\leq 12$.
@@ 2.3 Exact Results
The following is a corollary of Result 11 and Theorem 8.
@@
We have from Theorem 2 that ${{\pi}^{\sf c}}(P_{n})\geq{{\pi}^{*}}(P_{n})=\lceil\frac{2n}{3}\rceil$ and ${{\pi}^{\sf c}}(C_{n})\geq{{\pi}^{*}}(C_{n})=\lceil\frac{2n}{3}\rceil$.
We have from Theorem 14 that ${{\pi}^{\sf c}}(P_{n})\leq\lceil\frac{2n}{3}\rceil$. For $C_{n}$, partition $C_{n}$ into $\lfloor\frac{n}{3}\rfloor$ copies of $P_{3}$ and, possibly, an extra $P_{1}$ or $P_{2}$. Place two cops on the center vertex of each $P_{3}$, and one cop on each vertex of the remaining one or two vertices. The robber can only choose to start on one of the copies of $P_{3}$, where he is next to a pair of cops, and so will be captured on the first move. Thus ${{\pi}^{\sf c}}(C_{n})\leq\lceil\frac{2n}{3}\rceil$. ∎
@@
If $T$ is a tree with ${\sf rad}(T)=2$ and ${\sf diam}(T)=4$ then ${{\pi}^{\sf c}}(T)=4$.
@@
The upper bound follows from Result 4 and Theorem 12. The lower bound follows from Theorem 18 since $T$ contains $P_{5}$. ∎
Corollary 20.
If $T$ is a complete $k$-ary tree of depth 2 with $k\geq 3$, then ${{\pi}^{\sf c}}(T)=4$.
@@ 3 Cartesian Products
Let $G=P_{2}\Box P_{m}$ and for each $i\in[m]$ define the vertex subset $E_{i}=[2]\times\{i\}$. Suppose that $C$ can catch any robber on $G$. Then $C$ must be able to move two pebbles to any $E_{i}$. Indeed, assume that a cop catches the robber on $E_{i}$; without loss of generality, on vertex $(0,i)$. If the robber didn’t move, it is because each of its neighbors contained a cop. If the robber did move, say from $(1,i)$, then it was because a cop could have moved onto the robber. In each of those cases we see that two cops could be moved onto $E_{i}$. If the robber moved instead from, say, $(0,i-1)$ (or symmetrically $(1,i+1)$), then it was because cops could move into both $(0,i-1)$ and $(1,i-1)$ (or $(0,i+1)$ and $(1,i+1)$), and so two cops could be moved onto $E_{i-1}$ (respectively $E_{i+1})$. Now, by collapsing every $E_{i}$ we obtain the graph $G^{\prime}=P_{m}$, with corresponding collapsed configuration $C^{\prime}$. The above argument then shows by Result 19 that $C^{\prime}$ is 2-solvable. Consequently, Result 18 proves that $|C|=|C^{\prime}|\geq m+1$.
For the upper bound, we define $[k]=\{0,1,\ldots,k-1\}$, write $m$ uniquely as $m=4r+2s+t$, with $s,t\in[2]$, and let $V(G)=[2]\times[m]$. Next define the sets $S=\{(0,1)\}$ when $m$ is even and $S=\{(0,1),(m-1,s)\}$ when $m$ is odd, and $T=\{(4i+2j+1,j)\mid 0\leq i\leq r,j\in[2],4i+2j+1\in[m]\}$ — alternately, $T=\{(x,0)\mid x\equiv 1\pmod{m}\}\cup\{(x,1)\mid x\equiv 3\pmod{m}\}$. Then $S\cup T$ is a dominating set. Now place two cops on each vertex of $T$ and one cop on each vertex of $S$. It is simple to check that any robber can be captured in one step and that the number of cops in each case equals $m+1$. ∎
More general grids have cop pebbling numbers linear in their number of vertices, but there is a gap in the bounds for its coefficient.
@@
For all $16\leq k\leq m$ we have $\frac{5092}{28593}km+O(k+m)\leq{{\pi}^{\sf c}}(P_{k}\Box P_{m})\leq 2\left\lfloor\frac{(k+2)(m+2)}{5}\right\rfloor-8$. The lower bound also holds for all $1\leq k\leq m$.
@@
Result 21 and Theorem 2 produce the lower bound, while Result 12 and Theorem 5 produce the upper bound. ∎
For cubes, we can use Results 6 and 7 with Theorem 12 to obtain the upper bound ${{\pi}^{\sf c}}(Q^{d})\leq\lceil\frac{d+1}{2}\rceil d^{{\Theta}(d)}$. However, by adding extra cops in the Cops and Robbers game, we can reduce the capture time and therefore also reduce the upper bound for cop pebbling. Still, an exponential gap remains.
@@
The lower bound follows from Result 22 and Theorem 2. The upper bound follows from Result 13 and Theorem 5. ∎
@@
$\left(\frac{2}{3}\right)^{d}\sum_{u\in V(C_{k}^{d})}2^{-{\sf dist}(u,v)}$ $\leq\ \left(\frac{2}{3}\right)^{d}\sum_{i=0}^{kd/2}\binom{i+k-1}{k-1}2^{-i}\ \leq\ \left(\frac{2}{3}\right)^{d}\sum_{i=0}^{kd/2}\binom{i+k-1}{k-1}\ \leq\ \left(\frac{2}{3}\right)^{d}\binom{kd/2+k}{k}$
@@
by Fact 15 and Theorem 16. This, however, contradicts Theorem 2. ∎
@@
But we argue as follows that 13 pebble-cops on $(x,x)$ cannot catch a robber at $(v_{1},v_{1})$ in $G$. Compute all subscript arithmetic modulo $n-1$. We call a robber’s move from $(v_{i},v_{j})$ to $(v_{i+1},v_{j})$ (respectively $(v_{i-1},v_{j})$) increasing (respectively decreasing). First, it takes at least 4 pebble-cops to force the robber to move, because doing so requires 2 pebble-cops to be moved to a neighbor of the robber. Then, similarly, it takes at least 4 pebble-cops to prevent the robber from making successive increasing moves, and at least 4 more to prevent successive decreasing moves. Finally it requires at least 2 pebble-cops to block the robber’s movement inward, to some vertex $(x,v_{j})$. Note that all these sets of pebble-cops are distinct because no pebble-cop can perform two of the above actions simultaneously. But this is a contradiction because only 13 pebble-cops exist. Hence ${{\pi}^{\sf c}}(G)\geq 14$. ∎
@@ 4 Open Questions
Can the bounds .$4.178\approx\frac{5092}{28593}\leq\lim_{k,m\rightarrow\infty}{{\pi}^{\sf c}}(P_{k}\Box P_{m})/km\leq.$4 from Theorem 22 be improved?
Corollary 28 suggests the following two questions.
Question 30.
@@
Finally, Meyniel [13] conjectured in 1985 that every graph $G$ on $n$ vertices satisfies $c(G)=O(\sqrt{n})$. Some evidence in support of this conjecture is found in [4], where it is proved for $G\in{\cal G}_{n,p}$ that when $10<{\epsilon}<$ 1 and $p>2(1+{\epsilon})\log(n)/n$ we have $c(G)<\frac{10^{3}}{{\epsilon}^{3}}n^{\frac{1}{2}\log(n)}$ almost surely. (In fact, they also show that when $p\gg 1/n$ we have $c(G)>\frac{1}{(pn)^{2}}n^{\frac{1}{2}\left(\frac{\log\log(pn)-9}{\log\log(pn)}\right)}$ almost surely.) Along these lines, we make the following conjecture.
@@
Every graph $G$ on $n$ vertices satisfies ${{\pi}^{\sf c}}(G)\leq 2n/3+o(n)$; i.e., Ï(${}^{\sf c}(G)\geq n/3-o(n)$.

### v4

@@ 2.2 Upper Bounds
For integers $k$ and $d$, define the theta graph ${\Theta}(k,d)$ as the union of $k$ internally disjoint $xy$-paths, each of length $d$. Then ${\Theta}={\Theta}(k,2d)$ has n=k(d1)+2n=k(d-1)+2, $c({\Theta})=2$, ${\sf capt}_{2}({\Theta})=d$, ${\sf gir}({\Theta})=4d$, ${\gamma}({\Theta})=k\lceil(2d-3)/3\rceil$, and ${\gamma}_{d}=2$. Thus Theorem 5 yields an upper bound of roughly $4kd/3$, while Theorems 12 and 15 both yield the upper bound of $2^{d+1}$, which is better or worse than Theorem 5 when $k$ is bigger or less than, roughly, $3\cdot 2^{d-1}/d$.
@@
Consider the $(3,7)$-cage McGee graph $M$, defined by $V=\{v_{i}\mid i\in{\mathbb{Z}}_{24}\}$, with $v_{i}\sim v_{i+1}$ for all $i$, $v_{i}\sim v_{i+12}$ for all $i\equiv 0\pmod{3}$, and $v_{i}\sim v_{i+7}$ for all $i\equiv 1\pmod{3}$. We have ${\gamma}_{2}(M)\leq 4$ (e.g. $\{v_{0},v_{6},v_{9},v_{15}\}$), and so ${{\pi}^{*}}(M)\leq{{\pi}^{\sf c}}(M)\leq 16$ by Theorem 15. However, this bound is not tight, as ${{\pi}^{\sf c}}(M)\leq 12$: the vertex set $\{v_{i}\mid i\equiv 0\pmod{3}\}$ induces a matching of size 4 — for each edge, place 2 cops on one of its vertices and 1 cop on the other. Incidentally, this yields ${{\pi}^{*}}(M)\leq 12$; the best known lower bound on $\pi$$^{*}$} comes from Result 12: ${{\pi}^{*}}(M)\geq\lceil{\hat{\pi}^{*}}(M)\rceil=\lceil 64/7\rceil=10$. Hence we are left with a gap in the bounds for $M$: $10\leq{{\pi}^{*}}(M)\leq{{\pi}^{\sf c}}(M)\leq 12$.
@@ 2.3 Exact Results
The following is a corollary of Result 7 and Theorem 8.
@@
Result 10 and Theorem 2 provide the lower bound for both graphs. The upper bounds for paths and cycles follow from Theorems 14 and 3, respectively. ∎
@@
For any tree $T$ we have ${{\pi}^{\sf c}}(T)={{\pi}^{*}}(T)$.
@@
This follows from Theorems 2 and 4. ∎
A simple corollary to Theorem 19 is that any tree with radius 2 and diameter 4 has cop pebbling number 4. A complete $k$-ary tree of depth 2 with $k\geq 3$ is such an example.
@@ 3 Cartesian Products
Let $G=P_{2}\Box P_{m}$ and for each $i\in[m]$ define the vertex subset $E_{i}=[2]\times\{i\}$.
For the lower bound, we suppose that $C$ can catch any robber on $G$. Then we claim that $C$ must be able to move two pebbles to any $E_{i}$. Indeed, assume that a cop catches the robber on $E_{i}$; without loss of generality, on vertex $(0,i)$. (We shall assume throughout that the robber aims to maximize capture time.) If the robber didn’t move, it is because each of its neighbors contained a cop. If the robber did move, say from $(1,i)$, then it was because a cop could have moved onto the robber. In each of those cases we see that two cops could be moved onto $E_{i}$. If the robber moved instead from, say, (0,i1)(0,i-1) (or symmetrically $(1,i+1)$), then it was because cops could move into both (0,i1)(0,i-1) and (1,i1)(1,i-1) (or $(0,i+1)$ and $(1,i+1)$). At this point, the robber can’t move (else wouldn’t be captured at $(0,i)$), so there must be at least one cop at each of $(1,i)$ and $(0,i+1)$. The capture at $(0,i)$ implies that at least one of those vertices has two cops; in any case two cops can move to $E_{i}$. Now, by collapsing every $E_{i}$ we obtain the graph $G^{\prime}=P_{m}$, with corresponding collapsed configuration $C^{\prime}$. The above argument then shows by Result 14 that $C^{\prime}$ is 2-solvable. Consequently, Result 13 proves that $|C|=|C^{\prime}|\geq m+1$.
For the upper bound, we define $k]=\{0,1,\ldots,k-1\}$, write $m$ uniquely as $m=4r+2s+t$, with $s,t\in[2]$, and let $V(G)=[2]\times[m]$. Next define the sets $S=\{(0,1)\}$ when $m$ is even and $S=\{(0,1),(m-1,s)\}$ when $m$ is odd, and $T=\{(4i+2j+1,j)\mid 0\leq i\leq r,j\in[2],4i+2j+1\in[m]\}$ — alternately, $T=\{(x,0)\mid x\equiv 1\pmod{m}\}\cup\{(x,1)\mid x\equiv 3\pmod{m}\}$. Then $S\cup T$ is a dominating set. Now place two cops on each vertex of $T$ and one cop on each vertex of $S$. It is simple to check that any robber can be captured in one step and that the number of cops in each case equals $m+1$. (In fact, this is a Roman dominating set.) ∎
More general grids have cop pebbling numbers linear in their number of vertices, but there is a gap in the bounds for its coefficient. We use the following result on the Roman domination number of grids.
Result 17.
[13] For all all $k,m\geq 5$ we have ${\gamma}_{R}(P_{k}\Box P_{m})\leq\lfloor\frac{2(km+k+m)}{5}\rfloor$.
@@
For all $5\leq k\leq m$ we have $\frac{5092}{28593}km+O(k+m)\leq{\gamma}_{R}(P_{k}\Box P_{m})\leq\lfloor\frac{2(km+k+m)}{5}\rfloor$. The lower bound also holds for all $1\leq k\leq m$.
@@
Result 15 and Theorem 2 produce the lower bound, while Result 17 and Theorem 5 produce the upper bound. ∎
Note that this is another example of Roman domination giving a better upper bound than just domination (see Result 8).
For cubes, we can use Results 4 and 5 with Theorem 12 to obtain the upper bound ${{\pi}^{\sf c}}(Q^{d})\leq\lceil\frac{d+1}{2}\rceil d^{{\Theta}(d)}$. However, by adding extra cops in the Cops and Robbers game, we can reduce the capture time and therefore also reduce the upper bound for cop pebbling. Still, an exponential gap remains.
@@
The lower bound follows from Result 16 and Theorem 2. The upper bound follows from Result 9 and Theorem 5. ∎
@@
$\left(\frac{2}{3}\right)^{d}\sum_{u\in V(C_{k}^{d})}2^{-{\sf dist}(u,v)}\$$\leq\ \left(\frac{2}{3}\right)^{d}\sum_{i=0}^{kd/2}\binom{i+k-1}{k-1}2^{-i}\ \leq\ \left(\frac{2}{3}\right)^{d}\sum_{i=0}^{kd/2}\binom{i+k-1}{k-1}\ \leq\ \left(\frac{2}{3}\right)^{d}\binom{kd/2+k}{k}$
@@
by Fact 11 and Theorem 12. This, however, contradicts Theorem 2. ∎
@@
But we argue as follows that 13 pebble-cops on $(x,x)$ cannot catch a robber at $(v_{1},v_{1})$ in $G$. Compute all subscript arithmetic modulo n1n-1. We call a robber’s move from $(v_{i},v_{j})$ to $(v_{i+1},v_{j})$ (respectively $(v_{i-1},v_{j})$) increasing (respectively decreasing). First, it takes at least 4 pebble-cops to force the robber to move, because doing so requires 2 pebble-cops to be moved to a neighbor of the robber. Then, similarly, it takes at least 4 pebble-cops to prevent the robber from making successive increasing moves, and at least 4 more to prevent successive decreasing moves. Finally it requires at least 2 pebble-cops to block the robber’s movement inward, to some vertex $(x,v_{j})$. Note that all these sets of pebble-cops are distinct because no pebble-cop can perform two of the above actions simultaneously. But this is a contradiction because only 13 pebble-cops exist. Hence ${{\pi}^{\sf c}}(G)\geq 14$. ∎
@@ 4 Open Questions
Question 28.
Can the bounds $178\approx\frac{5092}{28593}\leq\lim_{k,m\rightarrow\infty}{{\pi}^{\sf c}}(P_{k}\Box P_{m})/km\leq.4$ from Theorem 21 be improved?
Corollary 27 suggests the following two questions.
@@
Finally, Meyniel [14] conjectured in 1985 that every graph $G$ on $n$ vertices satisfies $c(G)=O(\sqrt{n})$. Some evidence in support of this conjecture is found in [4], where it is proved for $G\in{\cal G}_{n,p}$ that when $0<{\epsilon}<1$ and $p>2(1+{\epsilon})\log(n)/n$ we have $c(G)<\frac{10^{3}}{{\epsilon}^{3}}n^{\frac{1}{2}\log(n)}$ almost surely. (In fact, they also show that when $p\gg 1/n$ we have $c(G)>\frac{1}{(pn)^{2}}n^{\frac{1}{2}\left(\frac{\log\log(pn)-9}{\log\log(pn)}\right)}$ almost surely.) Along these lines, we make the following conjecture.
@@
Every graph $G$ on $n$ vertices satisfies ${{\pi}^{\sf c}}(G)\leq 2n/3+o(n)$; i.e., $^{\sf c}(G)\geq n/3-o(n)$.
Data Availability Statement
No data were created or analyzed in this study.